# NeuralOS: Towards Simulating Operating Systems via Neural Generative Models

### v1

@@ 7 Experiments
Given the substantial computational resources required to train NeuralOS, our evaluation focused on NeuralOS variants, ablations, and intermediate training phases. For all evaluations, we used a subset of 730 examples from the reserved evaluation dataset. [^4] This number was chosen because our clustering procedure (detailed later) identified 73 clusters of challenging frame transitions, and we selected 10 examples per cluster, resulting in a total of 730 examples.
Figure 6: Ablation studies. (a) illustrates the limitations of directly using RNN outputs. (b) shows error accumulation without scheduled sampling, demonstrating its necessity for maintaining stability.
Cursor Position Accuracy
We evaluated cursor-position accuracy by training a regression model to predict cursor coordinates from the generated images. With the cursor position map, NeuralOS achieved highly accurate cursor localization, with an average position error of $\Delta x=1.6$ and $\Delta y=1.4$ pixels (Figure 5(b)). Given the $512\times 384$ resolution of the images, this corresponds to less than 0.5% of the frame width or height, indicating that cursor locations in generated images are very precise. This performance significantly outperformed a baseline without cursor position maps ($\Delta x=130.0$, $\Delta y=95.8$) [^5] This baseline is an earlier model version trained for 700K steps under slightly different conditions. and the random baseline ($\Delta x=175.4$, $\Delta y=126.9$), confirming the importance of explicit spatial encoding for accurate cursor localization.
@@
To evaluate state-transition modeling capability (e.g., opening applications), we clustered challenging frame transitions (those with mean pixel distance greater than 0.1 from the last input frame to the target frame, comprising approximately 2.8% of the dataset) into 73 categories. NeuralOS predictions, given identical past frames and user inputs, were matched against the closest cluster labels. As shown in Figure 5(a), NeuralOS achieved an accuracy of 37.7% (diagonal alignment), significantly outperforming majority voting (1.4%). We note that off-diagonal predictions may still represent valid outcomes due to inherent timing variability in OS actions (Appendix A).
Ablation Studies
@@
We further investigated the effectiveness of key training stages. Without the joint training stage (relying solely on the pretrained RNN), the predictions showed significant blurring (Figure 6(a)). This is caused by the MSE loss encouraging the RNN to predict averaged representations of multiple plausible outcomes rather than committing to a single clear target. Additionally, cursor positions were absent, despite the model correctly capturing state transitions (e.g., opening the home folder), indicating that the RNN still implicitly encoded cursor information.
@@
Furthermore, omitting the scheduled sampling stage led to rapid deterioration in generated frame quality due to compounding prediction errors over consecutive steps, as illustrated in Figure 6(b). In contrast, incorporating scheduled sampling greatly improved the model’s robustness (Figure 1).
8 Limitations
Our work represents an initial step toward a fully generative OS, but several significant limitations remain. Despite substantial training compute (17,000 H200 GPU hours and 6,000 H100 GPU hours), NeuralOS is still far from replicating the capabilities of a real operating system: screen resolution remains very low ($512\times 384$), fine-grained keyboard interactions such as accurately typing commands in a terminal are not reliably supported (see Appendix A for detailed failure cases), and inference is limited to approximately 1.8 fps using an NVIDIA H100 GPU. Additionally, many critical open challenges remain, including enabling the generative OS to install new software, interact with external resources (e.g., internet connectivity), and introduce controllability beyond traditional OS boundaries, similar to how language models can be steered by user intent (Hu et al., 2017). Addressing these limitations offers exciting avenues for future research that could realize the potential advantages of fully generative interfaces.
9 Related Work
NeuralOS is closely related to recent generative modeling approaches for simulating interactive environments conditioned on user inputs. “World Models”(Ha and Schmidhuber, 2018b) introduced latent-variable models for simulating reinforcement learning environments. GameGAN (Kim et al., 2020) used generative adversarial networks (GANs) for interactive game imitation, and Genie (Bruce et al., 2024) generated playable 2D platformer worlds. More recently, diffusion-based models have emerged as powerful real-time simulators: GameNGen (Valevski et al., 2024) simulated the game DOOM, MarioVGG (Protocol, 2024) simulated Super Mario Bros, DIAMOND (Alonso et al., 2024) simulated Atari and Counter-Strike, GameGen-X (Che et al., 2024) simulated open-world games, and Matrix (Feng et al., 2024) simulated AAA games. Beyond video games, UniSim (Yang et al., 2023) developed simulators for real-world scenarios, and Pandora (Xiang et al., 2024) introduced controllable video generation using natural-language prompts.
Compared to these prior works, NeuralOS addresses distinct challenges unique to OS simulation: while most GUI frame transitions involve subtle changes, accurately modeling critical discrete events, such as opening applications or menus, is essential. Additionally, precise cursor position prediction is crucial for interactive usability. NeuralOS introduces targeted model and training innovations specifically addressing these challenges, paving the way toward fully generative OS simulations.
@@ 10 Conclusion and Future Work
In this paper, we introduced NeuralOS, a neural framework for simulating operating system graphical interfaces through generative models. NeuralOS combines a recurrent neural network (RNN) for tracking persistent computer states with a diffusion-based neural renderer that generates screen images conditioned on user inputs. Trained on interaction examples from Ubuntu XFCE desktop environments, NeuralOS produces realistic screen sequences, accurately predicts mouse interactions, and captures state transitions such as opening applications. While precisely capturing fine-grained keyboard inputs remains challenging, NeuralOS provides a proof-of-concept demonstration of the feasibility of fully generative operating system interfaces.
NeuralOS opens several avenues for future research. First, future work could explore explicitly conditioning NeuralOS generations on natural-language user commands or gestures, enabling highly intuitive and fully customizable user experiences free from fixed interaction patterns. Second, NeuralOS’s neural architecture runs natively on parallel hardware such as GPUs, potentially enabling greater efficiency and richer interactions. Finally, fully generative operating systems may fundamentally blur the boundaries between traditionally distinct software categories, such as by converting passive media like movies into immersive video games directly at the OS level. NeuralOS is just an initial step towards a new computing paradigm, where user interactions and generative modeling converge to create intuitive, responsive, and adaptive systems. We hope NeuralOS can inspire a new generation of research at the intersection of generative modeling and interactive computing.
Acknowledgments and Disclosure of Funding
@@
This research was supported by Compute Canada through the Resources for Research Groups (RRG) 2025 competition, awarded to Yuntian Deng (RRG No. 5275), and was also partially supported by collaborative research funding from the National Research Council of Canada’s Artificial Intelligence for Design Program (AI4D-150). Additionally, Yuntian Deng acknowledges support from an NSERC Discovery Grant (RGPIN-2024-05178) and a Starter Grant provided by the University of Waterloo.
Appendix A Qualitative Analysis
Appendix B Full State Transition Heatmap
Appendix C Interactive Web Demo
To facilitate user interaction with NeuralOS, we developed a web-based frontend using FastAPI, accessible at https://neural-os.com/. Due to input rates (user actions) typically exceeding model inference speeds, we implemented a user-input queue. When the model finishes generating a frame, the system prioritizes processing recent meaningful inputs (clicks and keyboard events), discarding the redundant cursor movements if necessary. This approach maximizes the responsiveness of interactions.
Figure 10: Examples of original images (top row) and their corresponding reconstructions (bottom row) from the trained autoencoder. Despite significant spatial compression (8×8$\times$8 × downsampling), the autoencoder preserves details.
Appendix D Autoencoder Details
Model Architecture
Training
@@ Appendix D Autoencoder Details
Appendix E Multi-Stage Training Hyperparameters
This section provides detailed hyperparameters and dataset configurations for each training stage of NeuralOS, which is summarized in Table 1.
@@ Appendix E Multi-Stage Training Hyperparameters
Appendix F Computer-Use Agent Prompts

### v2

@@ 7 Experiments
Given the substantial computational resources required to train NeuralOS, our evaluation focused on NeuralOS variants, ablations, and intermediate training phases. For all evaluations, we used a subset of 730 examples from a reserved evaluation dataset, unless mentioned otherwise. [^1] This number was chosen because our clustering procedure (detailed later) identified 73 clusters of challenging frame transitions, and we selected 10 examples per cluster, resulting in a total of 730 examples.
Cursor Position Modeling
We evaluated cursor-position accuracy by training a regression model to detect cursor coordinates from the generated images. With the cursor position map, NeuralOS achieved highly accurate cursor localization, with an average position error of $\Delta x=1.6$ and $\Delta y=1.4$ pixels (Figure 4(b)). Given the $512\times 384$ resolution of the images, this corresponds to less than 0.5% of the frame width or height, indicating that cursor locations in generated images are precise. This performance significantly outperformed a baseline without cursor position maps [^2] This baseline is an earlier model version trained for 700K steps under slightly different conditions. ($\Delta x=130.0$, $\Delta y=95.8$) and the random baseline ($\Delta x=175.4$, $\Delta y=126.9$), confirming the importance of explicit spatial encoding for accurate cursor localization.
@@
Most OS interactions involve only cursor movements without significant visual changes. To focus evaluation on the crucial moments where the interface undergoes meaningful changes, such as opening or closing an application, we identified challenging frame transitions from the evaluation set. [^3] These were those with mean pixel distance greater than 0.1 from the previous frame to the target frame. These transitions were clustered into 73 categories using DBScan. NeuralOS predictions were then matched against the nearest cluster labels and compared with the ground truth. As shown in Figure 4(a), NeuralOS achieved an accuracy of 37.7% (correct state transitions over total state transitions) captured on the diagonal. This substantially outperformed majority voting (1.4%). Note that off-diagonal predictions may still correspond to valid outcomes due to inherent timing variability in OS actions (see Appendix D).
Figure 5: Doom interactions generated by NeuralOS. Doom was never installed in the underlying operating system used for data collection; instead, the model learned to simulate the application from synthesized training data, producing realistic walking and shooting behavior from user inputs.
Table 1: Human success rate (%) of identifying the real OS.
Crop Setting 10s 20s 30s 40s 50s 60s
No crop 58.3 55.0 60.0 56.7 61.7 59.3
Bottom cropped (40px) 51.7 53.3 50.0 48.3 46.7 50.0
Human Evaluation
We conducted a human evaluation to test how indistinguishable NeuralOS outputs are from a real operating system. Human raters were shown side-by-side video clips of the same user interaction sequence, one generated by NeuralOS and one recorded from an Ubuntu XFCE desktop, and asked to identify the real system. Each clip lasted 10–60 seconds (sampled at 10s intervals) and was drawn from real human interactions collected through our online demo.
To control for artifacts such as fluctuating disk-space counters (a known issue in diffusion), evaluation was performed under two conditions: (1) original videos and (2) videos with bottom 40 pixels (10% of height) cropped. In each condition, participants judged 30 randomly ordered comparisons.
As shown in Table 1, participants performed only slightly better than chance on up to 20 seconds of interactions, suggesting that NeuralOS is visually realistic for basic interactions.
Learning from Synthetic Demonstrations
A benefit of an end-to-end learned user interface is its ability to learn from demonstrations that do not correspond to real implementations. To illustrate this, we constructed a Doom application entirely from synthetic data, even though Doom was never installed in the docker. Desktop screenshots were augmented with a Doom icon, synthetic double-click actions were inserted to simulate launching the game, VizDoom gameplay segments with corresponding input events were spliced in, and each sequence ended with an ESC action returning to the desktop. Trained on these demonstrations, NeuralOS learned to launch, play, and close Doom as if it was a native application (Figures 1 and 5).
Long-Term Memory
We evaluated whether NeuralOS has long-term memory using a folder creation task. We navigated to Home $\to$ Documents, created a “New Folder” (via right-click $\to$ Create Folder), closed the file manager, and later reopened Documents after a delay. During training, the model saw only short delays: we included 500 examples where the file manager was reopened after 8 frames in the training set. At test time, we extended the delay to 64, 128, and 256 frames. As a control, we also included 500 examples where no folder was created.
Folder presence was evaluated by checking the region where the folder icon should appear, comparing it to the ground truth with an MSE threshold. Table 2 shows the model can mostly recall whether a folder had been created even after 256 frames, which is far beyond both the 8-frame training delay and its maximum training context of 64 frames. This indicates that NeuralOS develops persistent memory representations of application state that generalize well beyond its training horizon.
Figure 6: Ablation studies. (a) shows the limitations of directly using RNN outputs. (b) shows error accumulation without scheduled sampling, demonstrating its necessity for maintaining stability.
Table 2: Folder presence modeling confusion matrix after reopening file manager with increasing gaps. The higher the diagonal values (bolded), the better.
Frame Gap Ground Truth Predicted
Has Folder No Folder
64 Folder Created 0.52 0.48
No Folder Created 0.03 0.97
128 Folder Created 0.60 0.40
No Folder Created 0.00 1.00
256 Folder Created 0.62 0.38
No Folder Created 0.02 0.98
Ablations
@@
Without joint training (relying solely on the pretrained RNN), the predictions are blurry (Figure 6(a)). This is caused by the MSE loss encouraging the model to predict averaged representations of multiple plausible outcomes rather than committing to a single clear target. Additionally, cursor positions were absent, despite the model correctly capturing state transitions (e.g., opening the home folder), indicating that the RNN still implicitly encoded cursor information.
@@
Omitting scheduled sampling led to rapid deterioration in generated frame quality due to compounding prediction errors (Figure 6(b)). In contrast, incorporating scheduled sampling greatly improved the model’s robustness (Figure 1), which is also evident from quantitative evaluation in Figure 4(c).
@@ 8 Conclusion and Future Work
We introduced NeuralOS, a system that simulates OS graphical interfaces using generative models. Trained end-to-end on interaction sequences, NeuralOS produces realistic screen sequences, accurately predicts mouse interactions, and captures transitions such as opening applications.
A benefit of end-to-end trained interfaces is their ability to learn from synthetic demonstrations. Our Doom experiment illustrated this: the model simulated launching, playing, and closing Doom, even though the application was never installed in the underlying operating system. This points to a broader paradigm for future generative interfaces: if consistent demonstrations can be provided, even if they are manually edited, stylized, or fabricated, they can be internalized into usable interfaces.
Reproducibility Statement
Our training code and data processing scripts are available at https://github.com/yuntian-group/neural-os. An interactive demo is hosted at https://neural-os.com/, with demo code at https://github.com/yuntian-group/neuralos-demo. The model checkpoints used by the demo are publicly available on Hugging Face (linked directly from the demo repository). Additional details on training procedures and dataset construction are provided in the Appendix.
Ethics Statement
Generative models capable of imitating software inevitably raise broader considerations around intellectual property, licensing, and the economics of software distribution, similar to the concerns that emerged when large (vision) language models began reproducing copyrighted text, API patterns, or proprietary knowledge (He et al., 2025b; Wei et al., 2024; Cooper et al., 2025). To avoid these issues, training in this work was conducted on Ubuntu XFCE environments. More broadly, this project should be understood as an early research prototype investigating the feasibility of generative user interfaces rather than a system intended to replicate or redistribute commercial applications.
Acknowledgments
@@
This research was supported by Compute Canada through the Resources for Research Groups 2025 competition, awarded to Yuntian Deng (RRG No. 5275), and was also partially supported by collaborative research funding from the National Research Council of Canada’s Artificial Intelligence for Design Program (AI4D-150). Additionally, Yuntian Deng acknowledges support from an NSERC Discovery Grant (RGPIN-2024-05178) and a Starter Grant provided by the University of Waterloo.
Appendix A Baseline Comparison with DIAMOND
We adapted the DIAMOND model (Alonso et al., 2024), which was developed for action-conditioned video game simulation (e.g., Atari and CS:GO), to the operating-system setting. DIAMOND serves as a representative action-conditioned diffusion model capable of predicting future frames given past frames and actions.
Adaptation
DIAMOND was designed for video game environments with short-horizon temporal dependencies and coarse action spaces. To ensure a fair comparison, we modified it in the following ways:
  • Context window size: Increased from 4 $\to$ 64 past frames/actions to accommodate longer temporal dependencies in desktop interactions (e.g., launching Firefox may take 20–30 frames).
[…]
  • Keyboard action space: Expanded from the original 11 keys to the full keyboard (179 keys).
  • Training data: Trained on the same dataset used for NeuralOS to enable comparison.
Apart from these modifications, we followed the original architecture and training setup as closely as possible. Both DIAMOND and NeuralOS were trained from scratch under matched compute constraints (single B200 GPU for around 48 hours). NeuralOS in the main experiments was later trained longer (around 3M steps), but here we report results under equalized compute.
Qualitative Results
Figure 7 illustrates generations from the three models evaluated under the same interaction sequence used in the main paper (opening the “Home” folder, closing it, launching Doom, and returning to the desktop). The adapted DIAMOND model produces decent frames before the first abrupt transition occurs (opening Home), but generation quality collapses afterwards.
To improve the stability and reduce accumulated error over time, we also fine-tuned DIAMOND with scheduled sampling ($p=0.05$) for an additional 50k steps. This variant could recover from degenerations, but it still failed to execute application-launch transitions.
In contrast, the NeuralOS model produced more stable generations and was able to successfully open and close Doom, although at this training stage it had not yet learned to open the Home folder. (From the main experiments, we know that with longer training the model eventually learns this behavior.) Interestingly, the cursor is not rendered in these generations (we observed that learning to draw the cursor is particularly challenging because it occupies only a tiny fraction of the pixels and requires many gradient steps before it begins to appear). However, the model still responds to mouse hovers and double-clicks, indicating that the cursor position is being implicitly tracked even though its visual rendering has not yet emerged.
Quantitative Results
We additionally computed pixel RMSE between generated and ground-truth frames as a function of the number of generated steps. As shown in Figure 7, error for both DIAMOND variants increases rapidly, particularly after state-changing events such as opening an application. In contrast, NeuralOS maintains substantially lower error over longer horizons.
This comparison is not meant to imply that DIAMOND is unsuitable; rather, it highlights differences in modeling assumptions. Action-conditioned diffusion models developed for video games generally assume that the necessary state is visually encoded in recent frames, making short context windows sufficient. In operating-system settings, however, critical state may persist far beyond a short temporal window (e.g., whether a folder was created earlier in the workflow). With a fixed 64-frame context, DIAMOND cannot represent such long-term state by design. NeuralOS, through its recurrent state component, generalizes beyond its training horizon, consistent with the results in Table 2 showing correct folder-state recall after delays of up to 256 frames.
DIAMOND
DIAMOND + Scheduled Sampling ($p=0.05$)
NeuralOS
Figure 7: Generations from DIAMOND, DIAMOND with scheduled sampling, and NeuralOS evaluated on the same interaction trajectory used in Figure 1. Each pair of rows shows key frames as the model (a–c) opens and closes “Home”, followed by (d–f) launching and closing Doom.
Figure 8: Pixel RMSE of generated frames of NeuralOS, DIAMOND, and DIAMOND with scheduled sampling. Both DIAMOND variants show increasing error over frames, particularly after state transitions such as application launch events. NeuralOS maintains lower error across longer horizons, reflecting better state tracking enabled by its recurrent architecture.
$x_{t-2}$ $x_{t-1}$ $x_{t}$ $\hat{x}_{t}$
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 9: Correct prediction examples by NeuralOS. Each row shows two past frames (columns 1–2), ground-truth next frame (column 3), and NeuralOS’s prediction (column 4). Cursor positions are marked one frame in advance with circles (red: move-only, blue: left-click, yellow: right-click). NeuralOS correctly captures various GUI transitions, including opening menus and launching applications.
$x_{t-2}$ $x_{t-1}$ $x_{t}$ $\hat{x}_{t}$
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 10: Prediction examples where the generated frame does not match the ground truth frame. The layout follows Figure 9. Note that not all mismatches represent errors. For example, the third row illustrates a case where the screenshot tool window is closed in the ground truth frame but remains open in the prediction. This discrepancy arises because the window-closing action (not shown due to the limited context window) can have variable timing. Thus, both the predicted and ground truth frames are valid outcomes in such scenarios.
@@ Appendix A Baseline Comparison with DIAMOND
Table 4: Summary of the four-stage training pipeline. Each stage addresses a specific modeling challenge.
Stage Purpose Outcome Remaining limitations
Stage 1 – RNN Pretraining Train the RNN to predict latent frame embeddings so that it encodes cursor motion and keyboard inputs, ensuring the RNN is not ignored during later joint training. Model learns basic transitions; decoding the RNN output via the autoencoder yields plausible but blurry frames. Generations are blurry due to the MSE objective.
Stage 2 – Joint Training Combine the pretrained RNN with the diffusion-based renderer so that the decoder can sharpen predictions while retaining temporal structure from the RNN. Produces stable short-horizon interactions. Exposure bias: during training the model only sees ground truth frames, but at inference it must condition on its own imperfect outputs, leading to compounding errors over time.
Stage 3 – Scheduled Sampling Introduce scheduled sampling so the model is exposed to its own generated frames during training. Improves stability over extended interaction sequences. Context window length is still limited for training efficiency, and the model has not yet observed very long interaction horizons.
Stage 4 – Context Length Extension Expand the temporal context so the model can learn longer interaction dependencies. Enables the model to observe and learn longer-term dependencies. Future work is needed on efficiency, controllability, and external tool integration.
Appendix B Limitations
The current study uses Ubuntu XFCE as the operating environment. This choice reflects practical rather than technical constraints: Ubuntu provides a legally redistributable platform aligned with our goal of releasing all resources openly, and training comparable models across multiple operating systems would substantially increase data collection, storage, and compute requirements. Extending NeuralOS to more operating systems such as Windows or macOS remains an important direction for future work.
Our work represents an initial step toward a fully generative OS, but several limitations remain. Despite substantial training compute (17,000 H200 GPU hours and 6,000 H100 GPU hours), NeuralOS is still far from replicating the capabilities of a real operating system: screen resolution remains low, fine-grained keyboard interactions are not reliably supported (Appendix D). For inference, the current system requires an H100 GPU and is not yet suitable for cost-efficient or practical deployment. This reflects the exploratory nature of the work. We expect future architectures, training strategies, and hardware to make such systems substantially more efficient. Additionally, many open challenges remain, including enabling the generative OS to interact with external resources (e.g., internet), and introduce controllability beyond traditional OS boundaries.
Appendix C Multi-Stage Training Hyperparameters
NeuralOS is trained in four stages. A conceptual overview of the role and outcome of each stage is shown in Table 4, and the corresponding hyperparameter settings and dataset configurations are detailed in Table 3.
@@ Appendix C Multi-Stage Training Hyperparameters
Appendix D Qualitative Analysis
Appendix E Full State Transition Heatmap
Appendix F Interactive Web Demo
To make NeuralOS accessible, we built an interactive web demo available at https://neural-os.com. The system consists of a browser-based frontend and a GPU-backed backend.
Frontend
The frontend renders the NeuralOS screen inside a single HTML canvas. Mouse movements, clicks, and keyboard inputs are streamed to the backend over a persistent WebSocket connection.
Because user input often arrives faster than the model can produce frames (about 55 ms per frame), the system performs lightweight input coalescing: redundant mouse-move events are dropped when a new meaningful event (e.g., click or keyboard press) occurs. This makes interaction feel smoother.
Some OS actions require multiple generated frames to complete (for example, launching Firefox may take about 20 frames). To handle these cases, the demo includes an “Auto Input” mechanism: when the user stops interacting, the frontend periodically resends the last cursor position (every 67 ms) so that the model continues progressing through delayed transitions rather than waiting for new input.
Backend
The backend uses a dispatcher–worker architecture. The dispatcher manages queueing, session lifetime, WebSocket routing, and worker availability. Each worker process runs one instance of NeuralOS and occupies a single GPU. Only one user can actively interact per worker.
When users are waiting, the system applies time-limited sessions to ensure fair GPU access. The demo also includes an idle timeout, which triggers a “Connection Timeout” warning before disconnecting users who stop interacting.
To reduce bandwidth and improve responsiveness, generated images are split into $32\times 32$ pixel blocks, and only blocks that change meaningfully from the previous frame are transmitted to the browser.
Debug Controls
The demo includes internal debugging options such as toggling “Use RNN” (bypassing diffusion), changing sampling steps, resetting state, and modifying Auto Input behavior. These features were primarily used during development and are now hidden by default; advanced users can still access them via the browser console using showDebugControls().
Notably, “Use RNN” mode decodes the RNN state directly, which is only valid after the RNN pretraining stage. After joint training with the diffusion model, this mode produces degraded frames because it is no longer the intended decoding path. The default configuration (full RNN + diffusion model) corresponds to the model used in the paper.
Appendix G Transformer vs RNN Encoder
We initially implemented our system using a transformer-based encoder architecture. However, in interactive OS simulation, inference sequences can be arbitrarily long, as each user action generates a new frame that is appended to the model’s context unless explicitly truncated. For long or continuous interactions, this results in a steady increase in transformer memory consumption during inference, ultimately leading to out-of-memory (OOM) failures for sufficiently long sequences.
To address this, we replaced the transformer with a hierarchical recurrent neural network (RNN) architecture. RNNs maintain a constant-size hidden state between steps, enabling inference-time GPU memory usage that is independent of sequence length. Table 5 shows the memory requirements of a RNN encoder vs a Transformer encoder. Moreover, computer state modeling is inherently Markovian, making the RNN a suitable choice.
Table 5: Inference-time GPU memory usage (GB) for transformer and RNN architectures of similar parameter counts. The transformer encoder cannot accommodate 1024 frames on an L40 GPU.
Video Length (s) Transformer (GB) RNN (GB)
4 21.61 11.96
64 22.69 11.96
128 23.81 11.96
256 26.07 11.96
512 30.60 11.96
1024 >>40 (OOM) 11.96
Table 6: Application open–close accuracy. We report RMSE and mean absolute error (L1) between NeuralOS predictions and ground-truth frames, with pixel values normalized to [0, 1].
Application Action RMSE L1
Terminal Open 0.021 0.001
Firefox Open 0.048 0.006
Trash Open 0.017 0.001
Home Open 0.019 0.001
Terminal Close 0.015 0.001
Firefox Close 0.016 0.001
Trash Close 0.016 0.001
Home Close 0.016 0.001
Appendix H Application Open–Close Transitions
We evaluated the accuracy of application open–close transitions in NeuralOS. Specifically, we constructed 150 sequences for four main applications (Terminal, Firefox, Trash, Home) by sampling random cursor paths and click locations uniformly within the icon bounding boxes.
We compared the predicted and ground-truth frames at the end of each open and close event using root mean squared error (RMSE) and mean absolute error (L1). To account for rendering latency, we evaluated the frame after a fixed delay following each click: 40 frames after Firefox double-clicks, 10 frames after other application double-clicks, and 5 frames after single-click closing. These delays were determined empirically to align with the frame at which the application’s visual state change completes.
As shown in Table 6, the low RMSE and L1 scores confirm that NeuralOS accurately performs application open and close transitions, consistent with observations from the interactive demo.
Figure 12: Examples of original images (top row) and their corresponding reconstructions (bottom row) from the trained autoencoder. Despite significant spatial compression ($8\times$ downsampling), the autoencoder preserves details.
Appendix I Autoencoder Details
Model Architecture
Training
Appendix J Computer-Use Agent for Data Collection